OpenStudy (anonymous):

Let f(x) = x2 – 16. Find f^–1(x). please help! yes you have to inverse.

OpenStudy (mertsj):

x=y^2-16

OpenStudy (mertsj):

x+16=y^2 y=+/-sqrt(x+16)

OpenStudy (anonymous):

that is correct, thank you!(:

OpenStudy (mertsj):

yw